OpenStudy (anonymous):

Water coming out from a fountain is modeled by the function f(x) = -x2 + 5x + 4 where f(x) represents the height, in feet, of the water from the fountain at different times x, in seconds. What does the average rate of change of f(x) from x = 3 to x = 5 represent? The water falls down with an average speed of 3 feet per second from 3 seconds to 5 seconds. The water falls down with an average speed of 5 feet per second from 3 seconds to 5 seconds. The water travels an average distance of 3 feet from 3 seconds to 5 seconds. The water travels an average distance of 5 feet from 3 seconds

Vocaloid (vocaloid):

hint: x is the number of seconds average rate of change = speed

Vocaloid (vocaloid):

the next step is to calculate the average speed from x = 3 to x = 5 using the formula [f(5)-f(3)] / (5-3)

OpenStudy (anonymous):

So do i plug it in, into f(x) = -x2 + 5x + 4 or?????

Vocaloid (vocaloid):

yes

OpenStudy (anonymous):

so I do 5-3. Which is 2 so I plug 2 into everything?

Vocaloid (vocaloid):

no, not quite... f(5) - f(3) = ?

OpenStudy (anonymous):

f(2)

Vocaloid (vocaloid):

nope, not quite... f(x) = -x^2 + 5x + 4 f(5) = ?

OpenStudy (anonymous):

Im really confused?